What is the annual depreciation expense that we can take on our tax return based on the following data?

The SL method formula is – (Cost – salvage)/Life; Productivity unit method formula is ( (Cost – Salvage )/Estimated Life ) * # of miles driven in the year-

Cost of GM truck $55,000

Salvage value $5,000

Estimated Life 5 years or 150,000 miles

Drove the truck 80,000 miles in year 1.

Using straight-line method -What is our depreciation expense in year one? What is our book value – at the end of year-one? - Discuss

Using the production unit method - What is our depreciation expense in -year one? What is our book value – at the end of year-one?
Answer #1

Answers

  • Straight Line method

A

Cost

$            55,000.00

B

Residual Value

$              5,000.00

C=A - B

Depreciable base

$            50,000.00

D

Life [in years]

5

E=C/D

Annual SLM depreciation

$            10,000.00

>Depreciation expense in Year 1 = $ 10,000
>Book Value at the end of Year #1 = $ 55000 – 10000 = $ 45,000

  • Production Unit Method

A

Cost

$            55,000.00

B

Residual Value

$              5,000.00

C=A - B

Depreciable base

$            50,000.00

>Depreciation Expense in Year #1 = 80000 miles x ($50000 / 150000 miles) = $ 26,666.67 or $ 26,400(if rate rounded off)
>Book Value at the end of Year #1 = $ 55000 – 26400 = $ 28,600 or ($28333)